# CiteBART: Learning to Generate Citations for Local Citation Recommendation

### v2

@@ 1 Introduction
LCR has been addressed in a few works. BERT-GCN (Jeong et al., 2020) utilizes a feedforward neural network to combine local citation context representations using BERT with citation encodings through Graph Convolutional Neural Networks (GCN). The most recent solutions to the problem adopt a two-step process that consists of pre-fetching and re-ranking. DualEnh (Medić and Snajder, 2020) enhances a local citation context with the citing article’s title and abstract and uses this enhanced context as the query vector to retrieve the most similar candidate articles using their titles and abstracts. It performs this ranking through BiLSTM representations of inputs with attention layers on top. On the other hand, HAtten (Gu et al., 2022) initially pre-fetches a set of papers using the nearest neighbor search between local citation context extended with the citing paper’s title and abstract (query text as a whole) and the title and abstracts from a given pool of papers. Afterward, it re-ranks the selected candidate papers using a fine-tuned SciBERT (Beltagy et al., 2019) model where the input is the query text concatenated with a candidate paper’s title and abstract.
@@
Table 1: An example for input and target formats for pre-training and evaluation with CiteBART. Due to space constraints, we present the contexts and abstracts in an abbreviated form.
@@ 3.1 Custom BART Pre-training for LCR
We propose two training schemes for our approach: CiteBART-Base and CiteBART-Global. In CiteBART-Base, the model gets the masked context with the ground-truth citation as input. This setting tests the model’s performance in a local context-only situation (Table 1). With the underlying idea that good citation recommendation requires relating local citation contexts with the citing papers’ global information, such as titles and abstracts, we devised an innovative way to accomplish it. Inspiring from pre-training under the REALM framework (Guu et al., 2020), in CiteBART-Global, we append the citing paper’s title and abstract to the local context, allowing backpropagation through the global information that considers the pool of papers from the corpus. Specifically, we used the "$<$/s$>$" token designated by the pre-trained BART-base model as the separator.
@@ 3.2 Dataset Preprocessing
The datasets originally include a "TARGETCIT" marker as a placeholder for citations within each context. We replaced these markers with "$<$mask$>$" tokens to align with our pretraining process. Additionally, to ensure CiteBART focuses solely on predicting target citations, we removed any non-target citations from all four datasets.
@@ 4 Experiments
We conducted our experiments on devices with NVIDIA RTX6000 Ada GPU and NVIDIA V100 GPU [^6] Please find information on training and evaluation times in Appendix B.. The following hyperparameters were utilized in all our experiments. The number of epochs was set to $15$, as the change in loss values between epochs became negligibly small beyond this point. Only the PeerRead Global dataset has been trained for 30 epochs since the generative model requires longer training for the relatively smaller PeerRead dataset. We employed a learning rate of $52e-$ 5 and an attention dropout rate of $0.12$. Given that BART is a generative model, we adjusted its generation parameters to produce outputs that align with our requirements. Specifically, we utilized the grouped beam search with $20$ beams and applied a diversity penalty of $1.5$ to generate more diverse results. The maximum number of generated tokens was $25$ since the generated citations should not exceed it. Apart from these specific modifications, we did not alter the architecture of the BART model.
@@ 4.1 Results
We report our results using Recall@10 (R@10), Exact Match (EM), and Mean Reciprocal Rank (MRR) and compare with the state-of-the-art approaches in Table 4. As can be seen from the table, CiteBART-Global outperforms others on the existing benchmarks except for the smallest PeerRead dataset, while the base scheme is still a strong baseline, surpassing BERT-GCN on PeerRead, DualEnh, and HAtten on Refseer. The table includes the best-reported results of HAtten with $2k$ pre-fetched candidates. As for DualEnh (Medić and Snajder, 2020), we chose their superior "DualEnh-ws" model for the comparison. BERT-GCN’s (Jeong et al., 2020) results are available only on the PeerRead dataset. None of the past works have provided their Exact Match scores.
As shown in Table 4, CiteBART-Global demonstrates its advantage over HAtten on Refseer most since Refseer includes more training contexts compared to ArXiv. Given that CiteBART is a generative model, access to a larger training set contributes to its improved results.
Table 4: Comparison with state-of-the-art on LCR benchmarks.
@@
R@10 EM MRR R@10 EM MRR R@10 EM MRR R@10 EM MRR
BERT-GCNa - - - 0.529 - 0.418 - - - - - -
DualEnh-ws 0.703 - 0.366 - - - 0.534 - 0.280 - - -
HAtten 0.633 - -c 0.757 - -c 0.454b - -c 0.439b - -c
CiteBART-Base 0.686 0.422 0.504 0.570 0.363 0.424 0.606 0.382 0.449 0.355 0.184 0.240
CiteBART-Global 0.739 0.417 0.513 0.669 0.430 0.502 0.652 0.404 0.479 0.502 0.230 0.305
@@
  • b
    The reported results are based on a $10k$ subset of the test set.
  • c
    The authors did not report their MRR scores.
@@ 4.2 Qualitative Analysis
To provide insights into the working of CiteBART, we present some top $10$ prediction examples. We analyze four different scenarios shown in Table 5. Since CiteBART is a generative model, it is prone to hallucination. In the examples, the hallucinated predictions are designated with the * symbol. Furthermore, we provide another qualitative analysis on the performance of LLMs in LCR in Appendix C.
Table 5: Four example top-10 citation predictions using CiteBART. Due to space limitations, contexts and abstracts have been abbreviated. The hallucinated predictions are designated with the * symbol.
@@ 4.3 Ablation Study
Table 6: Ablation study results on ACL-200 Global dataset under four different configurations.
@@
The previous studies (Medić and Snajder (2020), Gu et al. (2022)) utilize an all-including training and inference configuration where citing and cited paper’s global information is concatenated with the local citation context. Their pre-fetch and re-ranking pipeline is well-suited to this setup and benefits from it as the inference step also allows incorporating the cited paper’s title and abstract, which is not the case in a learning approach like ours’. Ideally, CiteBART-Global outperforms them without exploiting global information about the cited papers.
@@ 4.4 Taxonomy and Measurement of Hallucinated Citations
In LCR, MaHR represents the proportion of hallucinated citations across all generated citations. As the task is evaluated with top-$k$ predictions for each test instance, the total number of responses becomes $k*n$ where $n$ is the number of test instances. Thus, MaHR is the fraction of hallucinated citations among $k*n$ responses (Equation 4).
@@ 4.5 Qualitative Analysis on Hallucinations
Table 9: Examples of hallucination categories. (a) No hallucination in any of the top-10 predictions. (b) Hallucinated publication years in the fourth, sixth, and ninth predictions. (c) Hallucinated author name in the sixth prediction. Fabricated author list in the ninth prediction. (d) Hallucinated author name in the fifth prediction. (A typo in the first author’s name). (e) Hallucinated author name in the sixth prediction (A single letter as the first author name). (f) In this cross-dataset example, CiteBART predicts a citation that has the same author name as the ground truth while in a different citation format and publication year.
@@ 5 Discussion and Conclusion
We comment on the advantages of using BART over encoder-based pre-training models such as RoBERTa. BART’s MLM objective is flexible and allows the masking of all the tokens in the parenthetical author-date style. RoBERTa cannot add citation tokens to its vocabulary by its MLM. Moreover, in the prediction phase, constraining predictions to citation tokens for RoBERTa is not straightforward. While BART is prone to hallucination, its capabilities significantly enhance LCR performance.
@@
For future work, we plan to investigate further the all-including configuration given in the ablation study. Conceptually, exploiting the cited paper’s title and abstract during the continual pre-training should have been complementary. However, the empirical evidence proves the contrary. More sophisticated masking strategies besides citation token masking should connect the dots by combining the information from the citing paper’s title and abstract, local citation context, and the cited paper’s title and abstract. Additionally, we should investigate the potential solutions to the citation-specific hallucinations and tackle a way to reduce the number of hallucinated recommendations in the top k.
@@ Acknowledgments
The Scientific and Technological Research Council of Türkiye (TUBITAK) supported this research with the 2219 fellowship awarded to the second author as a visiting scholar at the University of Edinburgh School of Informatics.
@@ Appendix A Token Limits
Before pre-training with citation objectives, we ensured that each context has its "$<$mask$>$" token in its middle position after tokenization. Another critical aspect was the determination of correct lengths for citation contexts. We limited citation contexts in each dataset to an optimal number of tokens to avoid increasing time and memory costs. An exploratory analysis of context lengths shows that the contexts of ACL-200 and Peerread are significantly longer than those of the other datasets. After tokenization, we observed that $200-400$ tokens were optimal for all base datasets. This limit allows sufficiently long contexts without a need for excessive amounts of padding tokens. As an exception, ACL-200 has $607$ contexts that exceed the $400$ limit. We have shortened them to the $400$ token limit as they correspond to a small proportion of the whole number of contexts and also because the number of discarded tokens is negligible.
@@ Appendix B Training and Evaluation Times
We conducted our experiments on devices with NVIDIA RTX6000 Ada GPU and NVIDIA V100 GPU for Global and Base datasets, respectively. For global datasets, the pre-training for Peerread and ACL-200 lasts for $2$ and $6$ hours, respectively. The larger datasets, Arxiv and Refseer, take up to $8-9$ days since they have similar sizes. For base datasets, the training for the smaller datasets, Peerread and ACL-200, lasts for 8 and 20 hours, respectively. The larger datasets, Arxiv and Refseer, take up to 14-15 days. However, we believe these relatively longer times are the result of training on the device with NVIDIA V100 GPU.
@@
Appendix C Qualitative Analysis on Large Language Models’ Performances in LCR
@@
We present four examples in Figures 1 and 2 to illustrate the workings of the base and global pre-training schemes, respectively. Due to space constraints, we partially display the list of citations, example contexts, and citing abstracts in the prompts. Each example consists of three parts: the prompt, the LLM’s answer, and the ground truth value of the masked citation token provided at the end of the prompt.

### v3

@@ 1 Introduction
Figure 1: CiteBART workflow. The yellow and green examples represent the workings of CiteBART-Base and CiteBART-Global, respectively. During inference, the expected outputs are in the author–date citation format, unlike the pre-training stage.
LCR has been addressed in a few works. BERT-GCN (Jeong et al., 2020) utilizes a feedforward neural network to combine local citation context representations using BERT with citation encodings through Graph Convolutional Neural Networks (GCN). The most recent solutions to the problem adopt a two-step process that consists of pre-fetching and re-ranking. DualEnh (Medić and Snajder, 2020) enhances a local citation context with the citing article’s title and abstract and uses this enhanced context as the query vector to retrieve the most similar candidate articles using their titles and abstracts. It performs this ranking through BiLSTM representations of inputs with attention layers on top. On the other hand, HAtten (Gu et al., 2022) initially pre-fetches a set of papers using the nearest neighbor search between local citation context extended with the citing paper’s title and abstract (query text as a whole) and the title and abstracts from a given pool of papers. Afterward, it re-ranks the selected candidate papers using a fine-tuned SciBERT (Beltagy et al., 2019) model where the input is the query text concatenated with a candidate paper’s title and abstract. SymTax (Goyal et al., 2024) improves upon HAtten by introducing an additional Enricher module and reranking candidate papers using taxonomical relationships along with contexts.
@@
Table 1: An example for input and target formats for evaluation with CiteBART. Due to space constraints, we present the contexts and abstracts in an abbreviated form.
@@
CiteBART presents a novel perspective to LCR. It achieves superior performance without relying on a pre-fetch and re-rank pipeline. It is an end-to-end learning system. Unlike previous works, we do not exploit the global information (titles and abstracts) of the target papers to make the recommendation. CiteBART-Global learns solely from the relation of citing papers’ global information with local citation contexts.
@@ 2 Related Work
SymTax (Goyal et al., 2024) introduces a three-stage recommendation architecture for the LCR task, consisting of the Prefetcher, Enricher, and Reranker modules. Prefetcher is the same as HAtten’s. Enricher leverages a pre-constructed citation network built from candidates to enhance their representation. Finally, Reranker combines a language model-based text relevance with a taxonomy relevance to yield a final recommendation. SymTax outperforms HAtten on the benchmark datasets.
@@ 3.1 Custom BART Pre-training for LCR
We propose two training schemes for our approach: CiteBART-Base and CiteBART-Global (Figure 1). In CiteBART-Base, the model gets the masked context with the ground-truth citation as input. This setting tests the model’s performance in a local context-only situation (Table 1). With the underlying idea that good citation recommendation requires relating local citation contexts with the citing papers’ global information, such as titles and abstracts, we devised an innovative way to accomplish it. Inspiring from pre-training under the REALM framework (Guu et al., 2020), in CiteBART-Global, we append the citing paper’s title and abstract to the local context, allowing backpropagation through the global information that considers the pool of papers from the corpus. Specifically, we used the "<</s>>" token designated by the pre-trained BART-base model as the separator.
@@ 3.2 Dataset Preprocessing
The datasets originally include a "TARGETCIT" marker as a placeholder for citations within each context. We replaced these markers with "<<mask>>" tokens to align with our pretraining process. Additionally, to ensure CiteBART focuses solely on predicting target citations, we removed any non-target citations from all four datasets.
@@ 4 Experiments
We conducted our experiments on devices with NVIDIA RTX6000 Ada GPU and NVIDIA V100 GPU [^6] Please find information on training and evaluation times in Appendix B.. The following hyperparameters were utilized in all our experiments. The number of epochs was set to $15$, as the change in loss values between epochs became negligibly small beyond this point. Only the PeerRead Global dataset has been trained for 30 epochs since the generative model requires longer training for the relatively smaller PeerRead dataset. We employed a learning rate of 2e52e-5 and an attention dropout rate of $0.12$. Given that BART is a generative model, we adjusted its generation parameters to produce outputs that align with our requirements. Specifically, we utilized the grouped beam search with $20$ beams and applied a diversity penalty of $1.5$ to generate more diverse results. The maximum number of generated tokens was $25$ since the generated citations should not exceed it. Apart from these specific modifications, we did not alter the architecture of the BART model.
@@ 4.1 Results
We report our results using Recall@10 (R@10) and Mean Reciprocal Rank (MRR) and compare with the state-of-the-art approaches in Table 4777We share our Exact Match (EM) scores in Appendix C.. As can be seen from the table, CiteBART-Global outperforms others on the existing benchmarks except for the smallest PeerRead dataset, while the base scheme is still a strong baseline.
HAtten reports its results based on a $10k$ subset of the test set due to long evaluation times. In Table 4, however, we present the results of HAtten on the entire test sets. As for DualEnh (Medić and Snajder, 2020), we chose their superior "DualEnh-ws" model for the comparison. BERT-GCN’s (Jeong et al., 2020) results are available only on the PeerRead dataset. We also compare our approach with SymTax (Goyal et al., 2024); its results surpass Hatten. Additionally, we add BM25 (Robertson and Zaragoza, 2009), a fast, TF-IDF-based retrieval function, as a baseline.
As shown in Table 4, CiteBART-Global demonstrates its advantage over SymTax and HAtten on Refseer most since Refseer includes more training contexts compared to ArXiv. Given that CiteBART is a generative model, access to a larger training set contributes to its improved results.
Table 4: Comparison with state-of-the-art on LCR benchmarks. The best values are shown with bold.
@@
R@10 MRR R@10 MRR R@10 MRR R@10 MRR
BM25 0.194 0.107 0.337 0.214 0.219 0.142 0.197 0.125
BERT-GCNa - - 0.529 0.418 - - - -
DualEnh-ws 0.703 0.366 - - 0.534 0.280 - -
HAtten 0.499 0.242 0.579 0.289 0.339 0.155 0.329 0.122
SymTax (SciV) 0.653 0.296 0.751 0.350 0.485 0.199 0.399 0.128
CiteBART-Base 0.686 0.504 0.570 0.424 0.606 0.449 0.355 0.240
CiteBART-Global 0.739 0.513 0.669 0.502 0.652 0.479 0.502 0.305
@@ 4.2 Qualitative Analysis
To provide insights into the working of CiteBART, we present some top $10$ prediction examples. We analyze four different scenarios shown in Table 5. Since CiteBART is a generative model, it is prone to hallucination. In the examples, the hallucinated predictions are designated with the * symbol.
Table 5: Four example top-10 citation predictions using CiteBART. Due to space limitations, contexts and abstracts have been abbreviated. The hallucinated predictions are designated with the * symbol. The correct predictions are in bold.
@@ 4.3 Ablation Study
Table 6: Ablation study results on ACL-200 Global dataset under four different configurations. The best values are shown with bold.
@@
The previous studies (Medić and Snajder (2020), Gu et al. (2022)) utilize an all-including training and inference configuration where citing and cited paper’s global information is concatenated with the local citation context. Their pre-fetch and re-ranking pipeline is well-suited to this setup and benefits from it as the inference step also allows incorporating the cited paper’s title and abstract, which is not the case in a learning approach like ours’. CiteBART-Global outperforms these models without relying on global information about the cited papers, representing a more ideal scenario for the LCR task.
@@ 4.4 Taxonomy and Measurement of Hallucinated Citations
In LCR, MaHR represents the proportion of hallucinated citations across all generated citations. As the task is evaluated with top-$k$ predictions for each test instance, the total number of responses becomes knk*n where $n$ is the number of test instances. Thus, MaHR is the fraction of hallucinated citations among knk*n responses (Equation 4).
@@ 4.5 Qualitative Analysis on Hallucinations
Table 9: Examples of hallucination categories. The referred predictions are in red. (a) No hallucination in any of the top-10 predictions. (b) Hallucinated publication years in the fourth, sixth, and ninth predictions. (c) Hallucinated author name in the sixth prediction. Fabricated author list in the ninth prediction. (d) Hallucinated author name in the fifth prediction. (A typo in the first author’s name). (e) Hallucinated author name in the sixth prediction (A single letter as the first author name). (f) CiteBART predicts a citation that has the same author name as the ground truth while in a different citation format and publication year. Unlike the other examples, the model’s pretraining dataset is different from the dataset associated with the given context.
@@ 10. Lin and Bilmes, 2008 *
4.6 LLMs in LCR
LLMs in LCR face a challenge retrieving the top $10$ citations for a given masked context. The main obstacle is the number of candidate citations in the citation pool, which contains $2043$ candidates, even for the smallest PeerRead. It is impractical for an LLM to evaluate every possible citation within a single prompt. Thus, the maximum context length and the size of the citation pool impose a significant bottleneck when applying LLMs to LCR.
To mitigate this issue, Jiang et al. (2025) proposed pre-fetching the top 100 candidates using a fast retrieval method such as BM25, and then passing only those candidates to the LLM prompt. Their experiments on the ArXiv and RefSeer datasets reported substantially lower $Recall@10$ scores ($0.134$ and $0.152$, respectively) than CiteBART. Their implementation presents each candidate in a separate prompt and asks for a similarity score in the range (01000-100) between the ground-truth and candidate citation to reach the overall ranking. As the approach requires $100$ separate prompts per example, the evaluation is prohibitively slow, and the produced similarity score in each case is not directly comparable to those of the others (many repetitive scores), lacking a sufficient basis for the final ranking.
Alternatively, we designed a prompt that simultaneously presented all $100$ pre-fetched citations and asked the LLM to select the top $10$. In practice, however, fitting citation metadata (titles and abstracts) into a single prompt often exceeded context length limits, and even when feasible, models frequently failed to select citations, producing invalid outputs. We also tested a simplified version, asking the LLM to return only the best citation for the exact match evaluation. Although this worked occasionally, the model often defaulted to echoing the top-ranked BM25 candidate. Our results suggest that the LCR task is currently quite challenging for LLMs due to prompt design and efficiency bottlenecks. We provide a qualitative analysis on the performance of LLMs in LCR in Appendix D.
@@ 5 Discussion and Conclusion
We comment on the pros of using BART over encoder-based pre-training models such as RoBERTa. BART’s MLM objective is flexible and allows the masking of all the tokens in the parenthetical author-date style. RoBERTa cannot add citation tokens to its vocabulary by its MLM. Moreover, constraining predictions to citation tokens for RoBERTa is not straightforward. While BART is prone to hallucination, its capabilities significantly enhance LCR performance.
@@
For future work, we plan to investigate further the all-including configuration given in the ablation study. Conceptually, exploiting the cited paper’s title and abstract during the continual pre-training should have been complementary. However, the empirical evidence proves the contrary. More sophisticated masking strategies besides citation token masking should connect the dots by combining the information from the citing paper’s title and abstract, local citation context, and the cited paper’s title and abstract. We also plan to investigate the connection between custom mask filling and the recognition of retrieval tokens in the context of generative information retrieval methods.We believe it is feasible to integrate custom citation mask-filling mechanisms with text generation models capable of producing citation placeholders. Additionally, we should investigate the potential solutions to the citation-specific hallucinations and tackle a way to reduce the number of hallucinated recommendations in the top k.
Limitations
We recognize the following limitations in this study. First, CiteBART addresses the task of LCR, predicting the best candidates for a citation placeholder in a given context. As a citation placeholder indicates that the context is worth citation, CiteBART builds upon the assumption of the citation worthiness of a local context.
Second, CiteBART necessitates pre-training on a specific dataset to recommend citations from the pool of papers in it. Thus, it may omit to cite some work or authors if they are not included in its training corpus. However, unlike the past works, as CiteBART is generative, it can recommend unseen papers, hallucinating. Although the fabricated citations in the top k predictions show that they capture the author names of the ground-truth citations, hallucination is still a problem.
Moreover, extending CiteBART to handle multi-citation scenarios, where a context refers to multiple citations simultaneously, would make the task setting more realistic for LCR. However, the current four LCR benchmarks only provide metadata (title and abstract) for the middle citation in each context, while other citations’ metadata are removed. Supporting multi-citation contexts would require minor modifications to our model architecture and codebase. Yet, more importantly, it necessitates constructing an LCR dataset specifically designed to include multiple citations (with all their metadata) per context.
There can be a bias towards citing papers as CiteBART learns from both local context and citing papers. Leveraging all the parts of a citation relationship, citing paper, local context, and cited paper should provide a more balanced learning process once it can be made learning. We leave this possibility for future exploration.
@@ Acknowledgments
The Scientific and Technological Research Council of Türkiye (TUBITAK) supported this research with the 2219 fellowship awarded to Selma Tekir as a visiting scholar at the University of Edinburgh School of Informatics. Selma is grateful to Mark Steedman for his hospitality and their fruitful discussions.
@@ Appendix A Token Limits
Before pre-training with citation objectives, we ensured that each context has its "<<mask>>" token in its middle position after tokenization. Another critical aspect was the determination of correct lengths for citation contexts. We limited citation contexts in each dataset to an optimal number of tokens to avoid increasing time and memory costs. An exploratory analysis of context lengths shows that the contexts of ACL-200 and Peerread are significantly longer than those of the other datasets. After tokenization, we observed that 200400200-400 tokens were optimal for all base datasets. This limit allows sufficiently long contexts without a need for excessive amounts of padding tokens. As an exception, ACL-200 has $607$ contexts that exceed the $400$ limit. We have shortened them to the $400$ token limit as they correspond to a small proportion of the whole number of contexts and also because the number of discarded tokens is negligible.
@@ Appendix B Training and Evaluation Times
We conducted our experiments on devices with NVIDIA RTX6000 Ada GPU and NVIDIA V100 GPU for Global and Base datasets, respectively. For global datasets, the pre-training for Peerread and ACL-200 lasts for $2$ and $6$ hours, respectively. The larger datasets, Arxiv and Refseer, take up to 898-9 days since they have similar sizes. For base datasets, the training for the smaller datasets, Peerread and ACL-200, lasts for 8 and 20 hours, respectively. The larger datasets, Arxiv and Refseer, take up to 14-15 days. However, we believe these relatively longer times are the result of training on the device with NVIDIA V100 GPU.
@@
Appendix C Exact Match Scores
Table 11 presents the exact match (EM) scores of CiteBART. While previous studies did not report EM scores, we consider this metric valuable for assessing the model’s ability to generate the correct citation on its first attempt. As shown in the table, CiteBART successfully predicts the correct citation directly for a substantial portion of the benchmark datasets.
Table 11: Exact Match (EM) score of CiteBART on LCR benchmarks.
Model ACL-200 PeerRead Refseer Arxiv
EM EM EM EM
CiteBART-Base 0.422 0.363 0.382 0.184
CiteBART-Global 0.417 0.430 0.404 0.230
Appendix D Qualitative Analysis on Large Language Models’ Performances in LCR
@@
We present four examples in Figures 2 and 3 to illustrate the workings of the base and global pre-training schemes, respectively. Due to space constraints, we partially display the list of citations, example contexts, and citing abstracts in the prompts. Each example consists of three parts: the prompt, the LLM’s answer, and the ground truth value of the masked citation token provided at the end of the prompt.